### v1

# Frontiers in Probabilistic Inference: Sampling Meets Learning

Abstract
In the field of machine learning coarse-grained potentials in molecular dynamics, propagators require that the Hamiltonian is quadratic in momentum, thus limiting the family of coarse-graining functions. In this paper, we derive a general family of coarse-graining embedding functions for which Langevin dynamics can be applied. This has significant implications in molecular simulations, and it paves the way for Langevin dynamics to be run on non-geometric coarse-graining representations such as those given by principal components of time-lagged independent component analysis (TICA) or latent embeddings of molecules obtained from neural networks.
1 Introduction
Coarse-grained molecular dynamics allows much faster simulations by reducing the number of degrees of freedom. The coarse grain function used determines which degrees of freedom are kept. We consider a fine-grained (FG) dynamical system as given by all-atom molecular dynamics. For example one approach in Husic et al. (2020) is only keeping the Cartesian coordinates of the Carbon $\alpha$ backbone of proteins. The FG system has microstates with positions and momentum $\vec{q},\vec{p}\in\mathbb{R}^{n}$ and Hamiltonian $H$. We aim to find a coarse-grained (CG) system that has macrostates with position $\vec{Q},\vec{P}\in\mathbb{R}^{N}$ with the mapping $\vec{Q}=f(\vec{q})$ for some latent space embedding function $f\in\mathbb{R}^{n}\rightarrow\mathbb{R}^{N}$. In the canonical ensemble, the constrained free energy for a macrostate as per Gibbs (1902) should satisfy:
$F(\vec{Q},\vec{P})=-k_{B}T\ln(Z(\vec{Q},\vec{P}))$
where $Z$ is the constrained partition function:
$Z(\vec{Q},\vec{P})=\frac{1}{h^{n}}\int_{V}d\vec{q}\int_{\mathbb{R}^{n}}d\vec{p}e^{-\frac{H(\vec{q},\vec{p})}{k_{B}T}}\prod_{i}^{N}\delta\Big(\vec{Q}_{i}-f_{i}(\vec{q})\Big)\prod_{i}^{N}\delta\Big(\frac{\partial F}{\partial P_{i}}-\sum_{j}^{n}\frac{\partial f_{i}(\vec{q})}{\partial q_{j}}\frac{\partial H}{\partial p_{j}}\Big)$ (1)
and $h$ is Plank’s constant. The first Dirac delta term ensures the positions of FG match the positions of CG macrostates, while the second term similarly constrains the velocities. Many bottom up CG interaction sites derivations such as Noid et al. (2008) constrain the embedding so the atoms $j\in I_{i}$ contributing weights $a_{ij}$ to interaction site $i$ are disjoint from all other interaction sites, so $I_{\alpha}\cap I_{\beta}=\emptyset$ if $\alpha\neq\beta$. Equivelently, it is a CG function that is just a straightforward linear transform $f(\vec{q})=\Xi\vec{q}$ where $\Xi\in\mathbb{R}^{N\times n}$, where $\Xi_{ij}=a_{ij}$, and the rows must be disjoint. In this case, assuming the Hamiltonian is of the form
$H(\vec{q},\vec{p})=\sum_{j}\frac{p_{j}^{2}}{2m_{j}}+U(\vec{q})$ (2)
the free energy, up to a constant, can be written in the form below.
$F(\vec{Q},\vec{P})=\sum_{i}\frac{P_{i}^{2}}{2M_{i}}+V(\vec{Q})$
where the effective CG masses are:
$M_{i}=\Big(\sum_{j\in I_{i}}\frac{a_{ij}^{2}}{m_{j}}\Big)^{-1}$ (3)
In this work, we claim that we can perform CG dynamics with a more general form of the Hamiltonian:
$H(\vec{q},\vec{p})=\frac{1}{2}\vec{p}^{\top}M^{-1}(\vec{q})\vec{p}+U(\vec{q})$ (4)
where $M$ is symmetric (if $M$ is not symmetric, symmetrizing it with $\frac{1}{2}(M+M^{\top})$ gives the same equations of motion). In addition, we show that for some choices of CG embedding function $f$, the free energy can be written in the form below:
$F(\vec{Q},\vec{P})=\frac{1}{2}\vec{P}^{\top}R^{-1}(\vec{Q})\vec{P}+V(\vec{Q})$ (5)
where $R(\vec{Q})$ is a function for the effective CG masses, satisfying
$R^{-1}(f(\vec{q}))=J_{f}(\vec{q})M^{-1}(\vec{q})J_{f}^{\top}(\vec{q})$ (6)
where $J_{f}$ is the Jacobian of $f$ and . In the previous case when $f(\vec{q})=\Xi\vec{q}$ with constant diagonal masses $M$, it reduces to $R^{-1}=\Xi M^{-1}\Xi^{\top}$ recovering 3. However, unlike previous methods requiring disjoint rows of $\Xi$, this method works for arbitrary $\Xi$ as long as $R$ is invertible. Here, $R$ must only depend on the codomain of $f$, such that $J_{f}M^{-1}J_{f}^{\top}$ is the same for all microstates $\vec{q}$ in the preimage of $\vec{Q}$, imposing a complex constraint on which $f$’s can be used. We show some general solutions and candidate $f$’s that satisfy these constraints in B.
2 Free Energy Potential and Momentum Contributions
In the case when the Hamiltonian is like in Eq. 4, substituting it in Eq. 1 and making some minor assumptions, we find the equation of the free energy to be
$e^{-\frac{F(\vec{Q},\vec{P})}{k_{B}T}}=\exp\Bigg(-\frac{1}{k_{B}T}\Big(\vec{P}^{\top}R^{-1}(\vec{Q})\vec{P}+V(\vec{Q})\Big)\Bigg)=\frac{O}{h^{n}}\int_{V}d\vec{q}\exp\Bigg(\\ -\frac{1}{k_{B}T}\bigg(U(\vec{q})-k_{B}T\ln\Big(\frac{C}{(2\pi)^{N}s^{N}O}\sqrt{\frac{(k_{B}T)^{-N}(2\pi)^{N}}{\text{det}\Big(J_{f}(\vec{q})M^{-1}(\vec{q})J_{f}^{\top}(\vec{q})\Big)}}\sqrt{\frac{(k_{B}T)^{n}(2\pi)^{n}}{\text{det}M^{-1}(\vec{q})}}\Big)\bigg)\\ -\frac{1}{k_{B}T}\Big(\vec{P}^{\top}(R^{-1}(\vec{Q}))^{\top}\Big(J_{f}(\vec{q})M^{-1}(\vec{q})J_{f}^{\top}(\vec{q})\Big)^{-1}R^{-1}(\vec{Q})\vec{P}\Big)\Bigg)\\ \prod_{i}^{N}\delta\Big(Q_{i}-f_{i}(\vec{q})\Big)$ (7)
where $C$ is the product of all the units of the $Q_{i}$’s, $O$ is the product of all the units of the $p$’s, and $s$ is the unit of time. One can see that if Eq. 6 is satisfied, then $V$ can be solved for as below:
$V(\vec{Q})=-k_{B}T\ln\Bigg(\frac{O}{h^{n}}\int_{V}d\vec{q}\exp\Bigg(-\frac{1}{k_{B}T}\bigg(U(\vec{q})\\ -k_{B}T\ln\Big(\frac{C}{(2\pi)^{N}s^{N}O}\sqrt{\frac{(k_{B}T)^{-N}(2\pi)^{N}}{\text{det}R^{-1}(\vec{Q})}}\sqrt{\frac{(k_{B}T)^{n}(2\pi)^{n}}{\text{det}M^{-1}(\vec{q})}}\Big)\bigg)\Bigg)\prod_{i}^{N}\delta\Big(Q_{i}-f_{i}(\vec{q})\Big)\Bigg)$ (8)
because the Dirac delta function removes the contributions of any points where the $R^{-1}(\vec{Q})$ doesn’t cancel in the last line of the exponent of 7, which can then be evaluated as a Gaussian.
3 Applications
3.1 Machine Learning
With machine-learned CG force fields, this method provides a way to separate the free energy into momentum and potential parts for a general set of CG functions $f$, rather than restricting them to a disjoint weighted sums of the Cartesian coordinates of atoms into interaction sites. Using this method, the neural network will only need to learn the potential portion of the free energy $V(\vec{Q})$ because the contribution of the momentum becomes trivial once one finds an expression for $R^{-1}(\vec{Q})$.
3.2 Langevin Dynamics
If the free energy is in the form of Eq. 5, it is known that Langevin dynamics can sample from a quadratic Hamiltonian of Eq. 4 with the SDE
$d\vec{q}=\nabla_{p}H(\vec{q},\vec{p})$
$d\vec{p}=\nabla_{q}H(\vec{q},\vec{p})-\gamma M^{-1}(\vec{q})\vec{p}+\sqrt{2\gamma k_{B}T}d\vec{W}$
where the steady state is the Boltzmann distribution as per Leimkuhler and Matthews (2015).
$\rho(\vec{q},\vec{p})\propto e^{-\frac{H(\vec{q},\vec{p})}{k_{B}T}}$
The macrostates can then also be sampled correctly with the SDE by substituting $H$ for $F$ and $M^{-1}(\vec{q})$ for $R^{-1}(\vec{Q})$ in the Langevin equation. The SDE will correctly sample the macrostates proportional to the free energy so that $\rho(\vec{Q},\vec{P})\propto e^{-\frac{F(\vec{Q},\vec{P})}{k_{B}T}}$. The Langevin equation will provide both “configurational” and “momentum” consistency (Jin et al. (2022)). The potential of mean force $\nabla_{Q}F(\vec{Q},\vec{P})$ (Ciccotti et al. (2005)) is expressed as
$\nabla_{Q}F(\vec{Q},\vec{P})=\Big\langle B(\vec{q})\nabla_{q}H(\vec{q},\vec{p})\Big\rangle-k_{B}T\Big\langle\nabla_{q}\cdot B(\vec{q})\Big\rangle$
and due to the quadratic separation can be written as
$\nabla_{Q}F(\vec{Q},\vec{P})=\Big\langle B(\vec{q})\nabla_{q}E(\vec{q})\Big\rangle_{f(\vec{q})=\vec{Q}}+\nabla_{Q}\Big(P^{\top}R^{-1}(\vec{Q})P\Big)-k_{B}T\Big\langle\nabla_{q}\cdot B(\vec{q})\Big\rangle_{f(\vec{q})=\vec{Q}}$
where $E$ is the exponent expression in Eq. 8, and $B(\vec{q})$ is any matrix satisfying the below.
$B(\vec{q})J^{\top}_{f}(\vec{q})=I$
One common choice is the pseudoinverse (if it exists) of the Jacobian.
Appendix A Proof for Equation 7
Proof.
Note that the derivation requires that $M$ is positive definite, and that $J_{f}M^{-1}J_{f}^{\top}$ is also positive definite. If $M$ is positive definite then $J_{f}M^{-1}J_{f}^{\top}$ will be positive definite as long as $J_{f}$ is full rank.
We start with 1 and substitute in the FG Hamiltonian.
$Z(\vec{Q},\vec{P})=\frac{1}{h^{n}}\int_{V}d\vec{q}\int_{\mathbb{R}^{n}}d\vec{p}e^{-\frac{\frac{1}{2}\vec{p}^{\top}M^{-1}(\vec{q})\vec{p}+U(\vec{q})}{k_{B}T}}\delta^{N}\Big(\vec{Q}-f(\vec{q})\Big)\prod_{i}^{N}\delta\Big(\frac{\partial F}{\partial P_{i}}-\sum_{j}^{n}\frac{\partial f_{i}(\vec{q})}{\partial q_{j}}\frac{\partial H}{\partial p_{j}}\Big)$
Move the terms only dependent on position out of the inner integral.
$\frac{1}{h^{n}}\int_{V}d\vec{q}e^{-\frac{U(\vec{q})}{k_{B}T}}\delta^{N}\Big(\vec{Q}-f(\vec{q})\Big)\int_{\mathbb{R}^{n}}d\vec{p}e^{-\frac{\frac{1}{2}\vec{p}^{\top}M^{-1}(\vec{q})\vec{p}}{k_{B}T}}\prod_{i}^{N}\delta\Big(\frac{\partial F}{\partial P_{i}}-\sum_{j}^{n}\frac{\partial f_{i}(\vec{q})}{\partial q_{j}}\frac{\partial H}{\partial p_{j}}\Big)$
Evaluate the expression inside the Dirac delta.
$\frac{1}{h^{n}}\int_{V}d\vec{q}e^{-\frac{U(\vec{q})}{k_{B}T}}\delta^{N}\Big(\vec{Q}-f(\vec{q})\Big)\int_{\mathbb{R}^{n}}d\vec{p}e^{-\frac{\frac{1}{2}\vec{p}^{\top}M^{-1}(\vec{q})\vec{p}}{k_{B}T}}\delta^{N}\Big(\nabla_{P}F(\vec{Q},\vec{P})-J_{f}(\vec{q})\nabla_{p}H(\vec{q},\vec{p})\Big)$
Assume $R^{-1}$ is symmetric.
$\frac{1}{h^{n}}\int_{V}d\vec{q}e^{-\frac{U(\vec{q})}{k_{B}T}}\delta^{N}\Big(\vec{Q}-f(\vec{q})\Big)\int_{\mathbb{R}^{n}}d\vec{p}e^{-\frac{\frac{1}{2}\vec{p}^{\top}M^{-1}(\vec{q})\vec{p}}{k_{B}T}}\delta^{N}\Big(R^{-1}(\vec{Q})\vec{P}-J_{f}(\vec{q})M^{-1}(\vec{q})\vec{p}\Big)$
Next, use the Fourier representation of the Dirac delta.
$\frac{1}{h^{n}}\int_{V}d\vec{q}e^{-\frac{U(\vec{q})}{k_{B}T}}\delta^{N}\Big(\vec{Q}-f(\vec{q})\Big)\int_{\mathbb{R}^{n}}d\vec{p}e^{-\frac{\frac{1}{2}\vec{p}^{\top}M^{-1}(\vec{q})\vec{p}}{k_{B}T}}\frac{s^{N}}{(2\pi)^{N}C}\int_{\mathbb{R}^{N}}d\vec{k}e^{i\vec{k}^{\top}\Big(R^{-1}(\vec{Q})\vec{P}-J_{f}(\vec{q})M^{-1}(\vec{q})\vec{p}\Big)}$
Assuming the necessary conditions, swap the integrals and move the non dependent terms out of the inner integral.
$\frac{s^{N}}{(2\pi)^{N}h^{n}C}\int_{V}d\vec{q}e^{-\frac{U(\vec{q})}{k_{B}T}}\delta^{N}\Big(\vec{Q}-f(\vec{q})\Big)\int_{\mathbb{R}^{N}}d\vec{k}e^{i\vec{k}^{\top}R^{-1}(\vec{Q})\vec{P}}\int_{\mathbb{R}^{n}}d\vec{p}e^{-\frac{\frac{1}{2}\vec{p}^{\top}M^{-1}(\vec{q})\vec{p}}{k_{B}T}-i\vec{k}^{\top}J_{f}(\vec{q})M^{-1}(\vec{q})\vec{p}}$
Complete the squares in the exponent.
$\frac{s^{N}}{(2\pi)^{N}h^{n}C}\int_{V}d\vec{q}e^{-\frac{U(\vec{q})}{k_{B}T}}\delta^{N}\Big(\vec{Q}-f(\vec{q})\Big)\int_{\mathbb{R}^{N}}d\vec{k}\\ e^{i\vec{k}^{\top}R^{-1}(\vec{Q})\vec{P}}\int_{\mathbb{R}^{n}}d\vec{p}e^{-\frac{1}{2}\Big(\vec{p}-ik_{B}TJ_{f}^{\top}k\Big)^{\top}\frac{M^{-1}(\vec{q})}{k_{B}T}\Big(\vec{p}-ik_{B}TJ_{f}^{\top}k\Big)-\frac{1}{2}k_{B}T\Big(J_{f}^{\top}\vec{k}\Big)^{\top}M^{-1}(\vec{q})\Big(J_{f}^{\top}\vec{k}\Big)}$ (9)
Move the nondependent part out.
$\frac{s^{N}}{(2\pi)^{N}h^{n}C}\int_{V}d\vec{q}e^{-\frac{U(\vec{q})}{k_{B}T}}\delta^{N}\Big(\vec{Q}-f(\vec{q})\Big)\int_{\mathbb{R}^{N}}d\vec{k}e^{i\vec{k}^{\top}R^{-1}(\vec{Q})\vec{P}-\frac{1}{2}k_{B}T\Big(J_{f}^{\top}\vec{k}\Big)^{\top}M^{-1}(\vec{q})\Big(J_{f}^{\top}\vec{k}\Big)}\\ \int_{\mathbb{R}^{n}}d\vec{p}e^{-\frac{1}{2}\Big(\vec{p}-ik_{B}TJ_{f}^{\top}k\Big)^{\top}\frac{M^{-1}(\vec{q})}{k_{B}T}\Big(\vec{p}-ik_{B}TJ_{f}^{\top}k\Big)}$ (10)
Since we assumed $M^{-1}$ is positive definite, then the right-most integral can be evaluated as a Gaussian integral. Gaussian integrals are invariant under translations, even if the translation is complex.
$\frac{s^{N}}{(2\pi)^{N}h^{n}C}\int_{V}d\vec{q}e^{-\frac{U(\vec{q})}{k_{B}T}}\delta^{N}\Big(\vec{Q}-f(\vec{q})\Big)\sqrt{\frac{(k_{B}T)^{n}(2\pi)^{n}}{\text{det}M^{-1}(\vec{q})}}\\ \int_{\mathbb{R}^{N}}d\vec{k}e^{i\vec{k}^{\top}R^{-1}(\vec{Q})\vec{P}-\frac{1}{2}k_{B}T\Big(J_{f}^{\top}\vec{k}\Big)^{\top}M^{-1}(\vec{q})\Big(J_{f}^{\top}\vec{k}\Big)}$ (11)
Complete the squares again.
$\frac{s^{N}}{(2\pi)^{N}h^{n}C}\int_{V}d\vec{q}e^{-\frac{U(\vec{q})}{k_{B}T}}\delta^{N}\Big(\vec{Q}-f(\vec{q})\Big)\sqrt{\frac{(k_{B}T)^{n}(2\pi)^{n}}{\text{det}M^{-1}(\vec{q})}}e^{-\frac{1}{k_{B}T}P^{\top}(R^{-1}(\vec{Q}))^{\top}\big(J_{f}(\vec{q})M^{-1}(\vec{q})J_{f}^{\top}(\vec{q})\big)^{-1}R^{-1}(\vec{Q})P}\int_{\mathbb{R}^{N}}d\vec{k}\\ e^{-\frac{1}{2}k_{B}T\big(\vec{k}+\frac{i}{k_{B}T}\big(J_{f}(\vec{q})M^{-1}(\vec{q})J_{f}^{\top}(\vec{q})\big)^{-1}R^{-1}(\vec{Q})P\big)^{\top}\big(J_{f}(\vec{q})M^{-1}(\vec{q})J_{f}^{\top}(\vec{q})\big)\big(\vec{k}+\frac{i}{k_{B}T}\big(J_{f}(\vec{q})M^{-1}(\vec{q})J_{f}^{\top}(\vec{q})\big)^{-1}R^{-1}(\vec{Q})P\big)}$ (12)
And once again assuming $\big(J_{f}(\vec{q})M^{-1}(\vec{q})J_{f}^{\top}(\vec{q})\big)^{-1}$ is positive definite, evaluate the Gaussian integral.
$\frac{s^{N}}{(2\pi)^{N}h^{n}C}\int_{V}d\vec{q}e^{-\frac{U(\vec{q})}{k_{B}T}}\delta^{N}\Big(\vec{Q}-f(\vec{q})\Big)\sqrt{\frac{(k_{B}T)^{n}(2\pi)^{n}}{\text{det}M^{-1}(\vec{q})}}\sqrt{\frac{(k_{B}T)^{-N}(2\pi)^{N}}{\text{det}\big(J_{f}(\vec{q})M^{-1}(\vec{q})J_{f}^{\top}(\vec{q})\big)}}\\ e^{-\frac{1}{k_{B}T}P^{\top}(R^{-1}(\vec{Q}))^{\top}\big(J_{f}(\vec{q})M^{-1}(\vec{q})J_{f}^{\top}(\vec{q})\big)^{-1}R^{-1}(\vec{Q})P}$ (13)
Appendix B Coarse Grain Function Candidates
B.1 Isometric Invariance
We assume that the standard case of $M$ is a position-independent diagonal matrix, of the same form as 2. Define the set of isometries that also commute with $M$ to be $G$ be all functions of $g(\vec{q})=R\vec{q}+b$ where $RMR^{\top}=M$ Here we will show that if $f$ is transitive on level sets, such that for every. If $f$ is transitive on level sets, so that $\forall\vec{q}_{1}\ \forall\vec{q}_{2}\ \ f(\vec{q}_{1})=f(\vec{q}_{2})\rightarrow\exists(g\in G)\ s.t.\ \vec{q_{1}}=g(\vec{q_{2}})$, and $f$ is invariant on the subset of $G$ used within level sets, so that $f=f\circ g$, then equation 6 can be satisfied. We can show that $J_{f}(\vec{q}_{1})MJ^{\top}_{f}(\vec{q}_{1})=J_{f}(\vec{q}_{2})MJ^{\top}_{f}(\vec{q}_{2})$ for any microstates $\vec{q}_{1},\vec{q}_{2}\in f^{-1}(\vec{Q})$ in the preimage of the same macrostate as follows:
Proof.
$J_{f}(\vec{q}_{1})MJ^{\top}_{f}(\vec{q}_{1})$
Existentially instantiate the $g$ for these two elements of the level set, and since $f$ is invariant on $g$
$J_{f\circ g}(\vec{q}_{1})MJ^{\top}_{f\circ g}(\vec{q}_{1})$
and the Jacobian of function composition is each Jacobian composed.
$J_{f}(g(\vec{q}_{1}))J_{g}(\vec{q}_{1})MJ_{g}^{\top}(\vec{q}_{1})J^{\top}_{f}(g(\vec{q}_{1}))$
$J_{f}(g(\vec{q}_{1}))RMR^{T}J^{\top}_{f}(g(\vec{q}_{1}))$
$J_{f}(\vec{q}_{2})MJ^{\top}_{f}(\vec{q}_{2})$
This proof shows the existence of an $R(Q)$ for any choices of $f$ that satisfies the conditions. ∎
B.2 Distance Latent Space
Consider a coarse grain function $f(\vec{q})$ that takes the pairs of distances between pairs of atoms. $f$ is isometric invariant almost by definition, so an $R$ should exist. We can find it by
$f_{ab}(\vec{q})=\sqrt{(q_{ax}-q_{bx})^{2}+(q_{ay}-q_{by})^{2}+(q_{az}-q_{bz})^{2}}$
Then
$(J_{f}MJ_{f}^{\top})_{ab,cd}=\sum_{i\in[n]}m_{i}\frac{\partial f_{ab}}{\partial\vec{q}_{i}}\frac{\partial f_{cd}}{\partial\vec{q}_{i}}$
Note that if $ab\cap cd$ do not share endpoints, then its just zero since one of the derivatives will always be zero. If they do, then let $b$ be the atom that they share.
$(J_{f}MJ_{f}^{\top})_{ab,bc}=\sum_{i\in[n]}m_{i}\frac{\partial f_{ab}}{\partial\vec{q}_{i}}\frac{\partial f_{bc}}{\partial\vec{q}_{i}}$
One can compute
$\frac{\partial f_{ab}}{\partial q_{ak}}=\frac{(q_{ak}-q_{bk})}{f_{ab}}$
If $ab=bc$ then
$(J_{f}MJ_{f}^{\top})_{ab,ab}=\sum_{k\in\{x,y,z\}}m_{a}\frac{\partial f_{ab}}{\partial\vec{q}_{ak}}\frac{\partial f_{ab}}{\partial\vec{q}_{ak}}+\sum_{k\in\{x,y,z\}}m_{b}\frac{\partial f_{ab}}{\partial\vec{q}_{bk}}\frac{\partial f_{ab}}{\partial\vec{q}_{bk}}=m_{a}+m_{b}$
$m_{a}\frac{f_{ab}^{2}}{f_{ab}^{2}}+m_{b}\frac{f_{bc}^{2}}{f_{bc}^{2}}=m_{a}+m_{b}$
otherwise
$(J_{f}MJ_{f}^{\top})_{ab,bc}=m_{b}\frac{\sum_{k\in\{x,y,z\}}\Big((q_{ak}-q_{bk})(q_{ck}-q_{bk})\Big)}{f_{ab}f_{bc}}=m_{b}\cos(\theta_{abc})$
Thus if $f$ provides enough edges so that the structure is "rigid", then the angles can be determined from the cosine rule and a formula for $R$ is trivial.